Gold, which has a density of 19.32 g/cm3, is the most ductile metal and can be pressed into a thin leaf or drawn out into a long fiber. (a) If a sample of gold with a mass of 2.075 g, is pressed into a leaf of 3.842 μm thickness, what is the area of the leaf? (b) If, instead, the gold is drawn out into a cylindrical fiber of radius 2.900 μm, what is the length of the fiber?
Answer #1

Part A.

Given that density of gold, rho = 19.32 gm/cm^3

Now given that gold leaf has, m = 2.075 gm

thickness of leaf, L = 3.842 \mu m = 3.842*10^-4 cm

Now we know that:

mass = density*Volume

Volume of leaf = Area*thickness

So,

m = rho*V = rho*A*L

A = m/(rho*L)

A = (2.075 gm/cm^3)/((19.32 gm/cm^3)*(3.842*10^-4 cm))

A = 2.075/(19.32*3.842*10^-4) cm^2

A = 279.55 cm^2 = 279.55*10^-4 m^2

A = Area of the leaf = 0.028 m^2

Part B.

Now when gold is drawn into cylindrical fiber of radius 2.900 \mu m, then

m = rho*A*L

A = cross-sectional area of cylinder = pi*r^2

L = m/(rho*A) = m/(rho*pi*r^2)

So,

L = 2.075 gm/((19.32 g/cm^3)*pi*(2.900*10^-4)^2 cm^2)

L = 2.075/(19.32*pi*(2.900*10^-4)^2) cm

L = 406504.26 cm = 406504.26*10^-2 m

L = 4065.0 m = length of fiber